OpenStudy (anonymous):

Getting stuck on integration by parts. Here's a similar problem to the one I'm trying to work out:

OpenStudy (anonymous):

\[\int\limits (x \cos(4x)) dx\]

OpenStudy (anonymous):

So if I understand correctly, I can look at this as two functions, where f(x)=cos(2x) and g'(x)=x.

OpenStudy (anonymous):

Ah, ok.

OpenStudy (loser66):

The integration by part : let u = x dv = cos (4x) dx du = dx v = \(\dfrac {sin(4x)}{4}\) the part u is quite easy, right? the part v : to get v, you just integral dv , and that is the result. Got this part?

OpenStudy (anonymous):

Yes.

OpenStudy (loser66):

then, just apply the formula

OpenStudy (anonymous):

now f(x)*g(x)-∫f'(x)g(x)?

OpenStudy (loser66):

yup

OpenStudy (anonymous):

Almost there...(coding all day, my brain is a little sloshy. :)

OpenStudy (loser66):

You can solve it, right? no stuck anymore, right?

OpenStudy (anonymous):

I thought so... So I take {[sin(2x)/2]*x}-∫[sin(2x)/2]dx

OpenStudy (loser66):

why 2x? the original problem is sin(4x) , right? why do you eat 2?

OpenStudy (anonymous):

Sorry, the original problem was 2x, I was trying to avoid the exact problem originally.

OpenStudy (loser66):

hahaha... big OK

OpenStudy (loser66):

next?

OpenStudy (anonymous):

I ended up with [(1/2)*x sin(x)]+[(1/4)cos(2x)]

OpenStudy (loser66):

+C

OpenStudy (anonymous):

Gah! I think I need to take a break! ;) 2x

OpenStudy (anonymous):

+ C, yes. :)

OpenStudy (anonymous):

That's the most this secular guy has ever wanted to shout "bless you!" Thanks for your help, I really appreciate it.

OpenStudy (loser66):

:)